A rocket of mass 14500kg accelerates at 210m/s2 for 32s from an initial speed of 5100m/s. (a) How fast will be rocket be travelling after the 32s?
(b) How much Kinetic Energy has the rocket gained?
1
Expert's answer
2019-07-19T14:38:19-0400

a)

"v=v_0+at=\\
= 5100+210\times 32= 11820 \frac{m}{s}"

b)

"E_{kin}=\frac{mv^2}{2}\\
\Delta E_{kin}=\frac{mv^2}{2}-\frac{mv_0^2}{2} = \\
\frac{14500\times 11820^2}{2}-\frac{14500\times 5100^2}{2} = 824 TJ"